OpenStudy (anonymous):

Prove whether the left expression is equal to the right expression and describe which property applies? 2(x − 3) 2x − 6

OpenStudy (anonymous):

it's equal..................... 2 * x =2x then 2* -3 = -6

OpenStudy (anonymous):

set them equal and try and get x = x for example prove if 2( x + 5) is equal to (2x + 10) 2(x + 5) = 2x + 10 2x + 10 = 2x + 10 x + 5 = x + 5 x = x